# Semantic Similarity Score for Measuring Visual Similarity at Semantic Level

### v1

@@ Abstract
Semantic communication, as a revolutionary communication architecture, is considered a promising novel communication paradigm. Unlike traditional symbol-based error-free communication systems, semantic-based visual communication systems extract, compress, transmit, and reconstruct images at the semantic level. However, widely used image similarity evaluation metrics, whether pixel-based MSE or PSNR or structure-based MS-SSIM, struggle to accurately measure the loss of semantic-level information of the source during system transmission. This presents challenges in evaluating the performance of visual semantic communication systems, especially when comparing them with traditional communication systems. To address this, we propose a semantic evaluation metric—SeSS (Semantic Similarity Score), based on Scene Graph Generation and graph matching, which shifts the similarity scores between images into semantic-level graph matching scores. Meanwhile, semantic similarity scores for tens of thousands of image pairs are manually annotated to fine-tune the hyperparameters in the graph matching algorithm, aligning the metric more closely with human semantic perception. The performance of the SeSS is tested on different datasets, including (1)images transmitted by traditional and semantic communication systems at different compression rates, (2)images transmitted by traditional and semantic communication systems at different signal-to-noise ratios, (3)images generated by large-scale model with different noise levels introduced, and (4)cases of images subjected to certain special transformations. The experiments demonstrate the effectiveness of SeSS, indicating that the metric can measure the semantic-level differences in semantic-level information of images and can be used for evaluation in visual semantic communication systems.
@@ B. Metric Variations under Different Compression Ratios
Figure 7 compares the SS, PSNR, and SSIM metrics of JPEG2000 at different bpp (bits per pixel) levels. It can be observed that as the compression ratio increases and the image details are continuously lost, the SS score also decreases accordingly, indicating that it has the capability to evaluate traditional image compression algorithms. In the horizontal comparison, the trends show that PSNR and SSIM are relatively even with respect to changes in bpp, but the changes in SS are more gradual in the high bpp region and decline rapidly in the low bpp region. This is consistent with human perception of images, as shown in the top right subfigure. When the compression rate is relatively low, the slight blurring and distortion in the image have a relatively small impact on the human’s semantic-level perception of the image. However, once a certain threshold is reached, the blurring and distortion make it difficult to recognize the content, and the human’s semantic-level perception of the image will decline rapidly.
Figure 8 shows the metric values of the JPEG image compression algorithm and the LSCI image semantic communication system[13] at different bpp (bits per pixel) levels, respectively. In each subfigure, the SS decreases as the bpp decreases, both in the traditional image compression algorithm and the deep learning-based semantic communication algorithm. This indicates that the SS metric can serve as a foundation for evaluating the performance of these algorithms. In contrast, other metrics such as MS-SSIM, LpipsScore, ViTScore, and ClipScore do not show obvious perception of image information loss at high bpp levels. In fact, metrics like ClipScore even exhibit the abnormal phenomenon of higher scores at lower bpp levels.
@@ C. Metric Variations under Different Signal-Noise-Ratio
Figure 13: Concrete examples and experiment statistic to illustrate the robustness of different metrics under changes weakly correlated with image semantics. Changes list here include small image translations, rotations, reflections, scaling, and color filter changes. SeSS is the most robust metric against these transformations.
@@ D. Metric Variations under Different Generation Noise
In contrast, ClipScore and SS are better able to measure the changes in the semantic-level information of the generated images, relative to the original images, under the influence of different noise strengths introduced by the generative model.
@@ E. Image Cases after Certain Special Transformations
Figure 13 uses visual examples to illustrate the robustness of different metrics under changes weakly correlated with image semantics, which list here include small image translations, rotations, reflections, scaling, and color filter changes. From a visual perspective, these changes have little impact on the semantic-level information in the images, and have a relatively small effect on human recognition of the semantic information. However, traditional pixel-level or structure-level metrics like PSNR or MS–SSIM give extremely low similarity scores, and patch-based metrics like LpipsScore and ViTScore also exhibit large fluctuations. The SS metric, on the other hand, demonstrates good robustness, indicating that it measures the difference in semantic-level information between images more effectively, making it more suitable for evaluating semantic-related tasks.
@@ V Conclusion
In this paper, a novel image similarity metric SeSS is proposed, which is based on the SAM segmentation model, Scene Graph Generation technique, and object-relation graph matching approach. The computation process of this metric exhibits structural and interpretable properties, enabling it to measure the semantic-level differences between images.
The semantic similarity of one hundred thousand image pairs has been manually annotated, and this annotated data has been utilized to optimize the hyperparameters of the graph matching algorithm within the SeSS, making the metric better aligned with human semantic perception. Subsequently, experiments were conducted to validate the performance of the SeSS.
@@
However, SeSS still has some limitations. Similar to the ViTScore, the SeSS values are often confined to the range above 0.3, failing to fully utilize the 0-1 interval. Furthermore, due to the internal algorithm architecture, the metric tends to perform better in evaluating the similarity of complex images containing multiple objects, while its accuracy may be inevitably lower for images without objects.
VI Acknowledgements

### v2

@@ Abstract
Semantic communication is considered a promising novel communication paradigm. Unlike traditional symbol-based error-free communication systems, semantic-based visual communication systems extract, transmit, and reconstruct images at the semantic level. However, widely used image similarity evaluation metrics, PSNR or MS-SSIM, struggle to accurately measure the loss of semantic-level information during transmission. This presents challenges in evaluating the performance of visual semantic communication systems, especially when comparing them with traditional systems. To address this, we propose a semantic evaluation metric—SeSS (Semantic Similarity Score), based on Scene Graph Generation and graph matching, which shifts the similarity scores between images into graph matching scores. Meanwhile, similarity scores for tens of thousands of image pairs are manually annotated to fine-tune the hyperparameters in the SeSS, aligning SeSS more closely with human semantic perception. The performance of the SeSS is tested on different datasets, including (1)traditional and semantic communication systems at different compression rates, (2)communication systems at different signal-to-noise ratios, (3)large-scale model with different noise levels introduced, and (4)cases of images subjected to certain special transformations. The experiments demonstrate the effectiveness of SeSS, indicating that SeSS can measure the semantic-level differences in semantic-level information of images and can be used for evaluation in visual semantic communication systems.
@@ B. Metric Variations under Different Compression Ratios
Figure 7 compares the SeSS, PSNR, and SSIM metrics of JPEG2000 at different bpp (bits per pixel) levels. It can be observed that as the compression ratio increases and the image details are continuously lost, the SeSS score also decreases accordingly, indicating that it has the capability to evaluate traditional image compression algorithms. In the horizontal comparison, the trends show that PSNR and SSIM are relatively even with respect to changes in bpp, but the changes in SeSS are more gradual in the high bpp region and decline rapidly in the low bpp region. This is consistent with human perception of images, as shown in the top right subfigure. When the compression rate is relatively low, the slight blurring and distortion in the image have a relatively small impact on the human’s semantic-level perception of the image. However, once a certain threshold is reached, the blurring and distortion make it difficult to recognize the content, and the human’s semantic-level perception of the image will decline rapidly.
Figure 8 shows the metric values of the JPEG image compression algorithm and the LSCI image semantic communication system[13] at different bpp (bits per pixel) levels, respectively. In each subfigure, the SeSS decreases as the bpp decreases, both in the traditional image compression algorithm and the deep learning-based semantic communication algorithm. This indicates that the SeSS metric can serve as a foundation for evaluating the performance of these algorithms. In contrast, other metrics such as MS-SSIM, LpipsScore, ViTScore, and ClipScore do not show obvious perception of image information loss at high bpp levels. In fact, metrics like ClipScore even exhibit the abnormal phenomenon of higher scores at lower bpp levels.
@@ C. Metric Variations under Different Signal-Noise-Ratio
Figure 13: Visual examples and experiment statistic to illustrate the robustness of different metrics under changes weakly correlated with image semantics. Changes list here include small image translations, rotations, reflections, scaling, and color filter changes. SeSS is the most robust metric against these transformations.
@@ D. Metric Variations under Different Generation Noise
In contrast, ClipScore and SeSS are better able to measure the changes in the semantic-level information of the generated images, relative to the original images, under the influence of different noise strengths introduced by the generative model.
@@ E. Image Cases after Certain Special Transformations
Figure 13 uses visual examples to illustrate the robustness of different metrics under changes weakly correlated with image semantics, which list here include small image translations, rotations, reflections, scaling, and color filter changes. From a visual perspective, these changes have little impact on the semantic-level information in the images, and have a relatively small effect on human recognition of the semantic information. However, traditional pixel-level or structure-level metrics like PSNR or MS–SSIM give extremely low similarity scores, and patch-based metrics like LpipsScore and ViTScore also exhibit large fluctuations. The SeSS metric, on the other hand, demonstrates good robustness, indicating that it measures the difference in semantic-level information between images more effectively, making it more suitable for evaluating semantic-related tasks.
@@ V Conclusion
In this paper, a novel image similarity metric SeSS is proposed, which is based on the SAM segmentation model, Scene Graph Generation technique, and object-relation graph matching approach. The computation process of this metric exhibits structural and interpretable properties, enabling it to measure the semantic-level differences between images. The similarity of image pairs has been manually annotated to optimize the hyperparameters of the graph matching algorithm within the SeSS, making the metric better aligned with human semantic perception. Subsequently, experiments were conducted to validate the performance of the SeSS.
@@
However, SeSS still has limitations. The SeSS values are often confined to the range above 0.3, failing to fully utilize the 0-1 interval. Furthermore, due to the internal algorithm architecture, the metric tends to perform better in evaluating the similarity of complex images containing multiple objects, while its accuracy may be inevitably lower for images without objects.
Figure 15: The similarity score of SeSS, ViTScore, PSNR, and MS-SSIM on the images compressed by JPEG, BPG, and the algorithm Text+Sketch at extreme low bpps. JPEG and BPG both reach their highest compression rates, with bpp at 0.090 and 0.038 respectively. The algorithm Text+Sketch compresses the images by extracting their natural language descriptions and Sketch features, and then generates the images using the Stable Diffusion model controlled by Controlnet, achieving a bpp of 0.015. According to the visual examples on the left, at extremely low compression rates, the JPEG and BPG compression algorithms exhibit significant distortion and blurriness, making the semantic information in the images difficult to recognize. In contrast, although Text+Sketch has differences in color and texture compared to the original image, it retains the main semantic information of the original image, and therefore receives a relatively higher SeSS score. However, on other metrics, the scores of the Text+Sketch algorithm are significantly lower than those of traditional compression algorithms.
Figure 16: Supplementary explanation for Experiment C. The Unclip model introduces different levels of noise to output images with varying degrees of similarity to the original image. From the visual examples on the left, when the noise is gradually increased to 600, the images’ semantic content compared to the original image changed slowly, and the difference in semantic levels gradually increases. However, when the noise is increased from 600 to 800, the semantic information undergoes a drastic change, leading to a significant difference between the generated image and the original image. The scores given by SeSS are consistent with this trend, demonstrating that SeSS is able to perceive the difference in semantic levels between the images generated by the image generation model and the original image.
This appendix mainly supplements the performance of SeSS when testing extremely compressed systems, the specific illustration of the test of image generated by Unclip model introducing different noises, and the specific performance of SeSS faced with changes weakly correlated with image semantics.
Figure 15 demonstrates the potential of SeSS to evaluate extremely compressed algorithms, as it is able to effectively focus on and assess the degree to which these algorithms preserve key semantic information during extreme compression. This encourages the development of more semantics-based extreme compression algorithms. Figure 16 provides visual examples that supplement the explanation of the reasonable trend of SeSS scores for Unclip-generated images under different noise levels. This proves that SeSS is able to perceive the difference in semantic levels between the images generated by the image generation model and the original image. Figure 17 further demonstrates the strong robustness of SeSS to changes weakly correlated with image semantics and its ability to perceive minor variations.
These results further demonstrate the advantages of SeSS in evaluating the outputs of image generation models, and its ability to provide objective evaluation scores in various domains, such as extreme compression. As a result, SeSS is expected to inspire more insightful work in the field of visual semantic communication.
Figure 17: Supplementary explanation for Experiment D. The similarity score of SeSS, MS-SSIM, Lpips_score, ViTScore, PSNR, on images changed by changes weakly correlated with image semantics. The images are changed with 4 different transformations: (1)Translation: The images are shifted to the left by varying numbers of pixels. (2)Rotation: The images are rotated counter-clockwise by varying degrees. (3) Scaling: The images are cropped from the top, bottom, left, and right by varying numbers of pixels, and then scaled back to the original size.(4) Color Filter: The images are overlaid with a pure blue image at varying blend ratios. SeSS demonstrates extremely strong robustness to these changes, and its metric values decreased smoothly as the degree of transformation increased.
Appendix A Acknowledgements